Pay Attention Please, it is a MATLAB Problem. Thanks!

2. f(t) 10u(t)-10u(t-10) +5u(t-10)-5u(t -20 Plot this function.
Answer #1

Ans)

Using a(t) = 1 for t > 0

a(t-r) = 1 for t >T

%==================MATLAB CODE for plot=========

t=0:0.1:30;

f=10*(t>=0)-10*(t>=10)+5*(t>=10)-5*(t>=20);

plot(t,f)

%================================

Plot obtained

LIVE EDITOR MEW E皐control ▼ 9。33- Aa Normal Run Section GoTo ▼ BIU M Run and Advance Find Text *-- Code- Refactor ▼ Section R

===================================